OpenStudy (anonymous):

the ratio of dimes to pennies in a purse if 3 to 4. if 3 pennies are taken out of the purse, the ratio of dimes to pennies becomes 1 to 1. how many dimes are in the purse

myininaya (myininaya):

d/p=3/4 d/(p-3)=1/1

myininaya (myininaya):

\[\frac{d}{p}=\frac{3}{4}\] \[4d=3p\] \[d=\frac{3p}{4}\] and we also have \[\frac{d}{p-3}=1\] \[d=p-3\] but d=3p/4 \[\frac{3p}{4}=p-3\] \[\frac{3p}{4}-p=-3\] \[\frac{3p}{4}-\frac{4p}{4}=-3\] \[\frac{-1}{4}p=-3\] \[p=-3(-4)=12\] so d=p-3=12-3=9

myininaya (myininaya):

9 dimes are in the purse